### 30 Dec 2023 v1

@@ 4.2 Practical Algorithms
$\hat{r}_{i}(s,a)=\frac{1}{N(s,a)}\sum_{\tau\in\mathcal{B}}\hat{\omega}(\tau|\mathcal{M}_{E})R_{j}(\tau).$ (15)
Here, $N(s,a)$ is the number of trajectories in the buffer $\mathcal{B}$ satisfying $f(s,a)\in f(\tau)$, and $\hat{\omega}(\tau|\mathcal{M}_{E})$ is also estimated by the MC method and is written as:
$\hat{\omega}(\tau|\mathcal{M}_{E}):=\frac{e^{-kd}}{\sum_{\tau\in\mathcal{B}}e^{-kd}+\epsilon},$ (16)
where $d=D_{\rm MMD}(\tau,\mathcal{M}_{E})$ is introduced in Section 3, $k$ and $\epsilon$ are predefined positive constants, and $\mathbb{I}_{\tau_{0}}{(\tau)}$ is the indicator function.
@@
For example, in a given 2D grid world shown in Fig. 1, the goal position is at the up-right of the starting location of the agent. A handful of previous good demonstration trajectories are provided with the agent ahead of time. Then, the proposed method can be used to approximate the value of the guidance reward with trajectories collected in each iteration. The estimation may be imprecise; however, we only ensure that lower guidance rewards are received by state-action pairs on trajectories far away from $\mathcal{M}_{E}$. In this manner, the agent is encouraged to approach the regions in which demonstration trajectories of $\mathcal{M}_{E}$ stretch. Therefore, our method avoids additional sampling for achieving stationary credit assignments and does not incur high computational costs compared to standard RL algorithms. Furthermore, we exploit past experiences gathered by the agent during training to update the previous good trajectory memory, which can help establish a more reliable guidance reward estimator.
@@ 5 Theoretical Analysis
Consider $\pi_{b}$ represents a behavior policy implied by the trajectory data set $\mathcal{P}$, and due to the optimality of trajectories in $\mathcal{P}$, it can be viewed as an expert policy with high returns. Therefore, we expect the training strategy to behave like $\pi_{b}$ and obtain a higher return. To achieve this goal, we propose to design a smooth guidance reward function $r_{i}(s,a)$ to assist policy optimization and help the agent policy gradually converge to $\pi_{b}$ under the current paradigm of RL. The following lemma illustrates that this guidance reward is an available reward function to ensure $\pi_{b}$ is the optimal policy. It is worth mentioning that, in implementation, the trajectories in $\mathcal{P}$ may not be optimal at the beginning of policy optimization, and $\mathcal{P}$ can be updated with the highly rewarded trajectories generated during the training process.
@@
Suppose $\pi_{b}$ is a policy implied by the replay memory $\mathcal{P}$ whose trajectories are optimal, and let the discount factor $\lambda$ be 1. According to the definition of the smooth reward $r_{i}(s,a)$, if the current policy $\pi$ is expressed as:
@@
$D_{\pi}(\pi^{\prime})\doteq\frac{T_{\pi}(\pi^{\prime})}{1-\gamma}-\frac{2\gamma\epsilon_{\pi^{\prime}}}{(1-\gamma)^{2}}\underset{s\sim d^{\pi}}{\mathbb{E}}\left[D_{TV}(\pi^{\prime}\parallel\pi)[s]\right],$
where $\gamma$ is the discount factor. $D_{TV}(\pi^{\prime}\parallel\pi)[s]=\frac{1}{2}\sum_{a}\left|\pi^{\prime}(a|s)-\pi(a|s)\right|$ is used to represent the total variational divergence between two action distributions of $\pi$ and $\pi^{\prime}$ when the state is $s$. The following bounds hold:
@@
Note that the bound in Theorem 1 we have given is based on TV divergence between policies. For ease of calculation, we further connect the performance bound to KL divergence through Pinsker’s inequality csiszar2011information : for two arbitrary distributions $p,q$, the TV and KL divergences satisfy the following in-equation: $D_{TV}(p\parallel q)\leq\sqrt{D_{KL}(p\parallel q)/2}$. Combining this with Jensen’s in-equation, we obtain
$\underset{s\sim d^{\pi}}{\mathbb{E}}\left[D_{TV}(\pi^{\prime}\parallel\pi)[s]\right]\leq\underset{s\sim d^{\pi}}{\mathbb{E}}\left[\sqrt{\frac{1}{2}D_{KL}(\pi^{\prime}\parallel\pi)[s]}\right]\leq\sqrt{\frac{1}{2}\underset{s\sim d^{\pi}}{\mathbb{E}}\left[D_{KL}(\pi^{\prime}\parallel\pi)[s]\right]}.$ (20)
@@
For any policies $\pi,\pi^{\prime}$ satisfying $\mathbb{E}_{s\sim d^{\pi}}[D_{KL}(\pi^{\prime}\parallel\pi)]\leq\delta$. Combining Eq. (20) with Eq. (19), we have:
@@ 6.1 Environments
Key-Door-Treasure Domain. We first evaluated the performance of POSG in the Key-Door-Treasure task whose state-action space is discrete, as shown in Fig. 1. The size of this grid-world environment is $26\times 36$. In each episode, the agent starts from the initial position in the bottom-left room of the Key-Door-Treasure domain. The maximum length of each episode is fixed, and an episode terminates immediately once the agent finds the treasure. The agent only receives a positive reward of 200 when it finds the treasure, and it cannot obtain any reward in other cases. At each time step, the agent is informed of its position information by the environment and chooses a possible action from the action space: move east, west, south, and north. To reach the location of the treasure, the agent is required to pick up the key (K) to open the door (D) and then travel through the room in the up-right corner to reach the treasure (T).
Figure 1: A collection of environments that we used to evaluate POSG: (a) Key-Door-Treasure domain; (b) SparseHalfCheetah; (c) SparseHopper; (d) Ant Maze.
Locomotion Tasks from MuJoCo. As shown in Figs. 1 and 1, we also demonstrated the effectiveness of POSG in several MuJoCo locomotion tasks with continuous state-action space. To investigate the potential limitations of POSG in more challenging learning tasks, we modified two classical MuJoCo agents, HalfCheetah and Hopper, and obtained two new agents named SparseHalfCheetah and SparseHopper: These agents yield a forward velocity reward only when the center-of-mass of the robot has already moved towards a certain direction for a threshold distance, and otherwise, the agent cannot obtain any positive reward. The threshold distance is 1 unit for SparseHopper and 10 units for SparseCheetah. At each time step, the agent observes the environment and performs an action sampled according to the policy. The agent receives an energy penalty caused by agent movement to adjust the torque applied to the robot joints.
Hard-Exploration Ant Maze. To further investigate the performance of POSG, we evaluated this algorithm with the ant maze task introduced as the benchmark for RL by duan2016benchmarking , which is depicted in Fig. 1. The hard-exploration property of this task is caused by two factors: continuous locomotion control and navigation in the maze. More specifically, the Ant robot must first learn to walk smoothly and then struggle to reach the target position in the maze. The Ant robot is only rewarded a large positive bonus when it reaches the specified position of the maze. Meanwhile, the episode will end when the agent receives the reward. The state space of this environment consists of two main parts: the agent’s internal joint angle information and task-specific attributes. The agent’s joint angle information is only determined by the agent’s internal state, and the task-specific attributes are mainly obtained from sensor readings, including the positions of walls and goals.
@@ 6.3 Baseline Methods
To investigate the benefits of using the smooth guidance rewards in the sparse reward setting, several different baseline methods are used for performance comparison in different tasks. More specifically, we carry out several state-of-the-art RL baselines, including SIL oh2018self , PPO+D libardi2021guided and generative adversarial self-imitation learning (GASIL) guo2018generative ; however, different from the standard SIL and GASIL framework in guo2018generative , our implemented GASIL is eligible to obtain the same demonstrations as POSG at the beginning of training. We also ran proximal policy optimization (PPO) schulman2017proximal baseline to verify the reward sparsity. For all baseline approaches, we adopt the parameters that produce the best performance during the parameter search. All performance curves were obtained by averaging over ten separate runs with different random seeds, and the shaded regions represent the standard error over these ten runs.
@@ 7.1 Evaluations in the Key-Door-Treasure domain
7.1.1 Performance Comparisons with Baseline Methods
As shown in Fig. 2, in the key-door-treasure domain, the sequential dependency property results in a low chance of completing all three tasks one by one: picking up the key, opening the door, and obtaining the treasure. We found that the PPO agent tends to get stuck at a sub-optimal policy that only picks up the key in a long time. It is worth noting that, for fairness, the SIL agent is provided with the same demonstration data as POSG at the beginning of policy optimization. The strong baseline, SIL and PPO+D, learn faster than PPO because these agents can obtain extra bonuses from the demonstration data. Specifically, SIL imitates and reproduces past good trajectories by storing experiences in the replay buffer and introducing a novel loss function. The PPO+D algorithm encourages agents to reach the goal quickly by replaying past good experiences. Interestingly, POSG can learn faster and is more stable than SIL, and POSG achieves competitive results against PPO+D in this sparse setting. This result indicates that our method can exploit the demonstrations better and obtain more useful and comprehensive information from these demonstrated trajectories since the smooth guidance reward design considers the demonstration’s sparse-reward return and distribution.
Figure 2: (a) Success rate in the Key-Door-Treasure domain; (b) The changing trend of the MMD distance; (c) State-action visitation graph of demonstrations; (d) State-action visitation graph of the POSG learned policy.
We further investigated the changing trend of the MMD distance for all methods during the training process. The changing curves of the MMD distance are depicted in Fig. 2. As shown in Fig. 2, the value of the MMD distance decreases to zero at the end of the training for the POSG, SIL, and PPO+D algorithms. This result indicates that these three approaches acquire the optimal behaviors that a demonstrator recommends. Furthermore, POSG’s MMD distance reduction rate is the highest among all methods. This result suggests that POSG avoids excessive meaningless exploration, and the smooth guidance reward provides more efficient policy optimization directions. Meanwhile, we draw the state-action visitation graphs of demonstrations and POSG learned policy in Figs. 2 and 2, respectively. The comparative result demonstrated that the POSG agent imitates experts’ behavior almost perfectly.
@@ 7.1.2 Visualization Comparisons for Exploration
Figure 3: The state-action visitation graphs of all algorithms: (a) PPO; (b) SIL; (c) PPO+D; (d) POSG.
@@ 7.2.1 Performance Comparisons with Baseline Methods
This experiment provided sparse-reward demonstrations with agents, each containing only one return scalar to suggest its merit. The dense OpenAI Gym reward signals were removed. We used the default OpenAI Gym reward function to measure the quality of actions learned with POSG and other baseline approaches. We compared the ground truth returns of different methods to illustrate the superiority of POSG.
Figs. 4 and 4 show that POSG outperforms other baseline approaches in the locomotion control tasks. Specifically, POSG can converge faster than different algorithms during policy optimization and achieve a higher final return. In contrast, PPO+D policy gradient computation only relies on the advantage information of demonstrations and thus cannot perform effective policy optimization with sparse-reward demonstrations. This causes the performance difference between PPO+D and POSG, and even the average return of PPO+D on the SparseHopper task no longer increases from the beginning of training. Since GASIL ignores reward signals of demonstrations and solely cares about the difference between the state-action visitation distributions of the current policy and the sparse-reward demonstration data, it cannot obtain competitive results compared to POSG. Consequently, this result emphasizes that POSG achieves simple and efficient credit assignments with sparse-reward demonstrations. POSG avoids information bottlenecks by integrating the sparse-reward demonstrations’ distribution information with the relevant trajectories’ returns.
Figure 4: (a) Learning curves of average return on the SparseHalfCheetah task; (b) Learning curves of average return on the SparseHopper task; (c) Learning curves of average ground truth reward per step on the SparseHopper task. The dotted line represents the return value of the demonstration trajectories under the ground truth reward function.
@@ 7.2.2 Evaluation of Ground-Truth Reward Learning
A deeper look into the experimental results is provided to verify further the POSG’s effectiveness in learning near-optimal behaviors. In this experiment, we implemented the PPO algorithm with ground-truth environmental rewards that are the default setting of OpenAI Gym openaigym . We then calculated the average default reward of each step. Meanwhile, we deployed the experiment with POSG and other baseline methods in Fig. 4 in the sparse reward version of the same task. The default OpenAI Gym’s reward function was used to measure the quality of actions learned with these approaches, and then the average ground-truth reward for each step was computed. We compared the learning curves of the average ground-truth rewards for different algorithms. The experimental results are shown in Fig. 4.
According to Fig. 4, the POSG average ground-truth reward dramatically increases at the beginning of learning, similar to PPO’s learning trend in the default OpenAI Gym reward setting. At the same time, its value achieves an incremental improvement in the process of policy optimization. Moreover, the average ground-truth reward of POSG gradually approximates that of PPO in the default reward setting at the end of training. This result indicates that the smooth guidance reward can help the agent learn near-optimal behaviors that are rewarded a high score by the default OpenAI Gym reward function. For a more intuitive comparison, Fig. 5 shows an example of two policies learned using POSG with sparse rewards and PPO with dense default rewards. Interestingly, the POSG agent has learned gaits similar to those of PPO. This fact further demonstrates the effectiveness of POSG in learning a near-optimal policy in tasks with sparse rewards.
Figure 5: (a) Agent trained by POSG; (b) Agent trained by PPO with the ground-truth reward function.

### 10 Apr 2024 v2

@@ 4.2 Practical Algorithms
$\hat{r}_{i}(s,a)=\frac{1}{N(x)}\sum_{\tau\in\mathcal{B}}\hat{\omega}(\tau|\mathcal{M}_{E})R_{j}(\tau)\mathbb{I}[x\in f(\tau)].$ (15)
Here, $x=f(s)$, $N(x)$ is the number of trajectories in the buffer $\mathcal{B}$ satisfying $x\in f(\tau)$, and $\mathbb{I}$ is the indicator function. $\hat{\omega}(\tau|\mathcal{M}_{E})$ is also estimated by the MC method and is written as:
$\hat{\omega}(\tau|\mathcal{M}_{E}):=\frac{e^{-kd(\tau)}}{\sum_{\upsilon\in\mathcal{B}}e^{-kd(\upsilon)}+\epsilon},$ (16)
where $d(\cdot)=D_{\rm MMD}(\cdot,\mathcal{M}_{E})$ is introduced in Section 3, $k$ and $\epsilon$ are predefined positive constants.
@@
For example, in a given 2D grid world shown in Fig. 1a, the goal position is at the up-right of the starting location of the agent. A handful of previous good demonstration trajectories are provided with the agent ahead of time. Then, the proposed method can be used to approximate the value of the guidance reward with trajectories collected in each iteration. The estimation may be imprecise; however, we only ensure that lower guidance rewards are received by state-action pairs on trajectories far away from $\mathcal{M}_{E}$. In this manner, the agent is encouraged to approach the regions where state-only demonstration trajectories of $\mathcal{M}_{E}$ stretch. Therefore, our method avoids additional sampling for achieving stationary credit assignments and does not incur high computational costs compared to standard RL algorithms. Furthermore, we exploit past experiences gathered by the agent during training to update the previous good trajectory memory, which can help establish a more reliable guidance reward estimator.
@@ 5 Theoretical Analysis
Consider that the trajectories in $\mathcal{M}_{E}$ are optimal, and $\pi_{b}$ represents a behavior policy implied by the trajectory data set $\mathcal{M}_{E}$. Then, due to the optimality of trajectories in $\mathcal{M}_{E}$, it can be viewed as an expert policy with high returns. Therefore, we expect the training strategy to behave like $\pi_{b}$ and obtain a higher return. To achieve this goal, we propose to design a smooth guidance reward function $r_{i}(s,a)$ to assist policy optimization and help the agent policy gradually converge to $\pi_{b}$ under the current paradigm of RL. The following lemma illustrates that this guidance reward is an available reward function to ensure $\pi_{b}$ is the optimal policy. It is worth mentioning that, in implementation, the trajectories in $\mathcal{M}_{E}$ may not be optimal at the beginning of policy optimization, and $\mathcal{M}_{E}$ can be updated with the highly rewarded trajectories generated during the training process.
@@
Suppose $\pi_{b}$ is a policy implied by the replay memory $\mathcal{M}_{E}$ which contains all optimal trajectories, and $p(s^{\prime}|s)=\pi_{b}(a|s)P(s^{\prime}|s,a)$ is the state transition function consistent with $\mathcal{M}_{E}$. Let the discount factor $\lambda$ be 1. According to the definition of the smooth reward $r_{i}(s,a)$, if the current policy $\pi$ is expressed as:
@@
$D_{\pi}(\pi^{\prime})\doteq\frac{T_{\pi}(\pi^{\prime})}{1-\gamma}-\frac{2\gamma\epsilon_{\pi^{\prime}}}{(1-\gamma)^{2}}\underset{s\sim d^{\pi}}{\mathbb{E}}\left[D_{\rm TV}(\pi^{\prime}\parallel\pi)[s]\right],$
where $\gamma$ is the discount factor. $D_{\rm TV}(\pi^{\prime}\parallel\pi)[s]=\frac{1}{2}\sum_{a}\left|\pi^{\prime}(a|s)-\pi(a|s)\right|$ is used to represent the total variational divergence between two action distributions of $\pi$ and $\pi^{\prime}$ when the state is $s$. The following bounds hold:
@@
Note that the bound in Theorem 1 we have given is based on TV divergence between policies. For ease of calculation, we further connect the performance bound to KL divergence through Pinsker’s inequality csiszar2011information : for two arbitrary distributions $p,q$, the TV and KL divergences satisfy the following in-equation: $D_{\rm TV}(p\parallel q)\leq\sqrt{D_{\rm KL}(p\parallel q)/2}$. Combining this with Jensen’s in-equation, we obtain
$\underset{s\sim d^{\pi}}{\mathbb{E}}\left[D_{\rm TV}(\pi^{\prime}\parallel\pi)[s]\right]\leq\underset{s\sim d^{\pi}}{\mathbb{E}}\left[\sqrt{\frac{1}{2}D_{\rm KL}(\pi^{\prime}\parallel\pi)[s]}\right]\leq\sqrt{\frac{1}{2}\underset{s\sim d^{\pi}}{\mathbb{E}}\left[D_{\rm KL}(\pi^{\prime}\parallel\pi)[s]\right]}.$ (20)
@@
For any policies $\pi,\pi^{\prime}$ satisfying $\mathbb{E}_{s\sim d^{\pi}}[D_{\rm KL}(\pi^{\prime}\parallel\pi)[s]]\leq\delta$. Combining Eq. (20) with Eq. (19), we have:
@@ 6.1 Environments
Key-Door-Treasure Domain. We first evaluated the performance of POSG in the Key-Door-Treasure task whose state-action space is discrete, as shown in Fig. 1a. The size of this grid-world environment is $26\times 36$. In each episode, the agent starts from the initial position in the bottom-left room of the Key-Door-Treasure domain. The maximum length of each episode is fixed, and an episode terminates immediately once the agent finds the treasure. The agent only receives a positive reward of 200 when it finds the treasure, and it cannot obtain any reward in other cases. At each time step, the agent is informed of its position information by the environment and chooses a possible action from the action space: move east, west, south, and north. To reach the location of the treasure, the agent is required to pick up the key (K) to open the door (D) and then travel through the room in the up-right corner to reach the treasure (T).
Fig. 1: A collection of environments that we used to evaluate POSG: (a) Key-Door-Treasure domain; (b) SparseHalfCheetah; (c) SparseHopper; (d) Ant Maze.
Locomotion Tasks from MuJoCo. As shown in Figs. 1b and 1c, we also demonstrated the effectiveness of POSG in several MuJoCo locomotion tasks with continuous state-action spaces. To investigate the potential limitations of POSG in more challenging learning tasks, we modified two classical MuJoCo agents, HalfCheetah and Hopper, and obtained two new agents named SparseHalfCheetah and SparseHopper: These agents yield a forward velocity reward only when the center-of-mass of the robot has already moved towards a certain direction for a threshold distance, and otherwise, the agent cannot obtain any positive reward. The threshold distance is 1 unit for SparseHopper and 10 units for SparseCheetah. At each time step, the agent observes the environment and performs an action sampled according to the policy. The agent receives an energy penalty caused by agent movement to adjust the torque applied to the robot joints.
Hard-Exploration Ant Maze. To further investigate the performance of POSG, we evaluated this algorithm with the ant maze task introduced as the benchmark for RL by duan2016benchmarking , which is depicted in Fig. 1d. The hard-exploration property of this task is caused by two factors: continuous locomotion control and navigation in the maze. More specifically, the Ant robot must first learn to walk smoothly and then struggle to reach the target position in the maze. The Ant robot is only rewarded a large positive bonus when it reaches the specified position of the maze. Meanwhile, the episode will end when the agent receives the reward. The state space of this environment consists of two main parts: the agent’s internal joint angle information and task-specific attributes. The agent’s joint angle information is only determined by the agent’s internal state, and the task-specific attributes are mainly obtained from sensor readings, including the positions of walls and goals.
@@ 6.3 Baseline Methods
To investigate the benefits of using the smooth guidance rewards in the sparse reward setting, several different baseline methods were used for performance comparison in different tasks. We first compared our method with the state-of-the-art RL methods that learn from state-only demonstrations. To achieve this goal, we adopted GAIfO torabi2018generative as the baseline. However, this method is a state-only imitation learning method, and it does not interact with the environment to gain new experiences during training. For a fair comparison, we combined GAIfO with SIL oh2018self and obtained a new RL method GASIfO. Furthermore, we used several other state-of-the-art RL baselines, including SIL, PPO+D libardi2021guided and generative adversarial self-imitation learning (GASIL) guo2018generative . Unlike the standard SIL and GASIL framework in guo2018generative , our implemented GASIL could obtain the same demonstrations as POSG at the beginning of training. It is worth that these three methods can access the demonstrator’s action information. We also ran proximal policy optimization (PPO) schulman2017proximal baseline to verify the reward sparsity. For all baseline approaches, we adopted the parameters that produced the best performance during the parameter search. All performance curves were obtained by averaging over ten separate runs with different random seeds, and the shaded regions represented the standard error over these ten runs.
6.4 Acquisition of offline state-only demonstrations
We used the PPO algorithm based on the default OpenAI Gym reward function to train expert policies for each task, shown in the figures as expert. Then, we used these policies to generate expert-level trajectories and only stored the observation and return information. We selected and saved policies learned during training as the medium-level policies named medium in the corresponding figures. We used these policies to generate medium-level trajectories and only preserved the observation and return information.
@@ 7.1 Evaluations in the Key-Door-Treasure domain
7.1.1 Performance Comparison with Baseline Methods
As shown in Fig. 2a, in the key-door-treasure domain, the sequential dependency property results in a low chance of completing all three tasks one by one: picking up the key, opening the door, and obtaining the treasure. We found that the PPO agent tends to get stuck at a sub-optimal policy that only picks up the key in a long time. It is worth noting that, for fairness, the SIL agent is provided with the same demonstration data as POSG at the beginning of policy optimization. The strong baselines, SIL and PPO+D, learn faster than PPO because these agents can obtain extra bonuses from the demonstration data. Specifically, SIL imitates and reproduces past good trajectories by storing experiences in the replay buffer and introducing a novel loss function. The PPO+D algorithm encourages agents to reach the goal quickly by replaying past good experiences. Interestingly, POSG can learn faster and is more stable than SIL, and POSG achieves competitive results against PPO+D in this sparse setting. This result indicates that our method can exploit the demonstrations better and obtain more useful and comprehensive information from these demonstrated trajectories since the smooth guidance reward design considers the demonstration’s sparse-reward return and distribution.
Fig. 2: (a) Success rate in the Key-Door-Treasure domain; (b) The changing trend of the MMD distance; (c) State-action visitation graph of demonstrations; (d) State-action visitation graph of the POSG learned policy.
We further investigated the changing trend of the MMD distance for all methods during the training process. The changing curves of the MMD distance are depicted in Fig. 2b. As shown in Fig. 2b, the value of the MMD distance decreases to zero at the end of the training for the POSG, SIL, and PPO+D algorithms. This result indicates that these three approaches acquire the optimal behaviors that a demonstrator recommends. Furthermore, POSG’s MMD distance reduction rate is the highest among all methods. This result suggests that POSG avoids excessive meaningless exploration, and the smooth guidance reward provides more efficient policy optimization directions. Meanwhile, we draw the state-action visitation graphs of demonstrations and POSG learned policy in Figs. 2c and 2d, respectively. The comparative result demonstrated that the POSG agent imitates experts’ behavior almost perfectly.
@@ 7.1.2 Visualization Comparisons for Exploration
Fig. 3: The state-action visitation graphs of all algorithms: (a) PPO; (b) SIL; (c) PPO+D; (d) POSG.
@@ 7.2.1 Performance Comparisons with Baseline Methods
This experiment provided sparse-reward demonstrations with agents, and each trajectory contained only one return scalar to suggest its merit. The dense OpenAI Gym reward signals were removed. We used the default OpenAI Gym reward function to measure the quality of actions learned with POSG and other baseline approaches and computed the ground truth returns with this default OpenAI Gym reward function. We compared the ground truth returns of different methods to illustrate the superiority of POSG.
Figs. 4a and 4b show that POSG outperforms other baseline approaches in the locomotion control tasks. Specifically, POSG can converge faster than different algorithms during policy optimization and achieve a higher final return. In contrast, PPO+D policy gradient computation only relies on the advantage information of demonstrations and thus cannot perform effective policy optimization with sparse-reward demonstrations. This causes the performance difference between PPO+D and POSG, and even the average return of PPO+D on the SparseHopper task no longer increases from the beginning of training. Since GASIL and GASIfO ignore reward signals of demonstrations and solely care about the distribution difference between the current policy and the sparse-reward demonstration data, they cannot obtain competitive results compared to POSG. The performance difference between GASIL and GASIfO in Fig. 4a demonstrates the importance of action information for learning. These results show that POSG achieves simple and efficient credit assignments with sparse-reward demonstrations. POSG avoids information bottlenecks by integrating the sparse-reward demonstrations’ distribution information with the relevant trajectories’ returns.
Fig. 4: (a) Learning curves of average return on the SparseHalfCheetah task; (b) Learning curves of average return on the SparseHopper task; (c) Learning curves of average ground truth reward per step on the SparseHopper task. The dotted line represents the return value of the demonstration trajectories under the ground truth reward function.
@@ 7.2.2 Evaluation of Ground-Truth Reward Learning
A deeper look into the experimental results is provided to further verify the POSG’s effectiveness in learning near-optimal behaviors. In this experiment, for comparison, we trained the PPO algorithm with ground-truth environmental rewards that are the default setting of OpenAI Gym openaigym . We then calculated the average default reward of each step. Meanwhile, we deployed the experiment with POSG and other baseline methods in Fig. 4b in the sparse reward version of the same task. The default OpenAI Gym’s reward function was used to measure the quality of actions learned with these approaches, and then the average ground-truth reward for each step was computed. We compared the learning curves of the average ground-truth rewards for different algorithms. The experimental results are shown in Fig. 4c.
According to Fig. 4c, the POSG average ground-truth reward dramatically increases at the beginning of learning, similar to PPO’s learning trend in the default OpenAI Gym reward setting. At the same time, its value achieves an incremental improvement in the process of policy optimization. Moreover, the average ground-truth reward of POSG gradually approximates that of PPO in the default reward setting at the end of training. This result indicates that the smooth guidance reward can help the agent learn near-optimal behaviors that are rewarded a high score by the default OpenAI Gym reward function. For a more intuitive comparison, Fig. 5 shows an example of two policies learned using POSG with sparse rewards and PPO with dense default rewards. Interestingly, the POSG agent has learned gaits similar to those of PPO. This fact further demonstrates the effectiveness of POSG in learning a near-optimal policy in tasks with sparse rewards.
Fig. 5: (a) Agent trained by POSG; (b) Agent trained by PPO with the ground-truth reward function.